OpenStudy (anonymous):

please help me in diff calculus! :)) can you show the proper solving of this one? f(x) = 1/2x^2 + 3x 1/x^2? diff calculus

OpenStudy (anonymous):

\[f(x)=\frac{1}{2}x^2+3x+\frac{1}{x^2}\] or something else?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

ahh no

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

yeah thats it soory

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

ahhm can you help to solve that one?

OpenStudy (anonymous):

yeah ahhm very sure

OpenStudy (anonymous):

i would be happy to unfortunately you do not have a question, just a function what is the question? since you wrote "calc" it could be a) graph b) find the derivative c) identify the critical points d) find the anti derivative

OpenStudy (anonymous):

i am kind of hoping it is 'find the derivative" because that would be easiest

OpenStudy (anonymous):

my prof said we dont need to get the grapgh just derivatives only

OpenStudy (anonymous):

so the question is given \(f(x)=\frac{1}{2}x^2+3x+\frac{1}{x^2}\) find \(f'(x)\) right?

OpenStudy (anonymous):

thats why its so hard to me :(

OpenStudy (anonymous):

yeah f ^1(x)

OpenStudy (anonymous):

repeated applications of the all mighty power rule, most easily written as \[\frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

ok ill get my reading glass

OpenStudy (anonymous):

for example the derivative of \(x^7\) is \(7x^6\) and the derivative of \(x^{-3}\) is \(-3x^{-4}\)

OpenStudy (anonymous):

\[\huge \frac{d}{dx}[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

than how can we solve if we get the proper formula?

OpenStudy (anonymous):

i am not sure what you are asking your job is to find the derivative right? and the tool we have at our disposal is the "power rule" so for example the derivative of \[\huge \frac{1}{2}x^2\]is \[\huge 2\times \frac{1}{2}x^{2-1}=x\]

OpenStudy (anonymous):

the derivative of \(3x\) is just \(3\) since the derivative of a line is the slope

OpenStudy (anonymous):

the only one that requires some work is to find the derivative of \(\frac{1}{x^2}\) to use the power rule write \[\frac{1}{x^2}=x^{-2}\] and then use it

OpenStudy (anonymous):

@satellite73 show me the proper, its hard to me

OpenStudy (anonymous):

show me the proper way of salving, to easy to understand sir.

OpenStudy (anonymous):

we take the derivative piece by piece \[f(x)=\frac{1}{2}x^2+3x+x^{-2}\] \[f'(x)=2\times \frac{1}{2}x^{2-1}+3x^{1-1}-2x^{-2-1}\] \[f'(x)=x+3-2x^{-3}\]

OpenStudy (anonymous):

so the final answr is x+3 -2x^2?

OpenStudy (anonymous):

no but close final ansswer is \[x+3-2x^{-3}\] or \[x+3-\frac{2}{x^3}\]

OpenStudy (anonymous):

@satellite73 thank you sir .:)

OpenStudy (anonymous):

yw hope it helped